OpenStudy (kaylala):

How many grams is 1 a.m.u. ? (please provide or show a Detailed SOLutioN) Thanks!

thomaster (thomaster):

An atomic mass unit is equal to 1.66053892 × 10^-27 kilograms

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20090107072739AA1Pe1D hope it helps ...

OpenStudy (kaylala):

@thomaster so, how do you compute for the grams?

thomaster (thomaster):

What do you mean? what 1 amu is in gram? That's \(\large\sf 1.66053892 × 10^{-24}\) gram

OpenStudy (kaylala):

@thomaster what is 1 amu in gram? --- this is what i meant now, how did you get 1.66 x 10^-24 gram? i need the solution too. if you may

thomaster (thomaster):

It is just a constant, you can't calculate it. A long time ago they decided to take the mass of carbon-12 and divide that by 12 to get 1 amu.

OpenStudy (kaylala):

@thomaster but how come my book has a derivation / solution of how to know how much is 1g in amu? see the attached picture below:

OpenStudy (kaylala):

HERE IS THE SOL'n

thomaster (thomaster):

that's the relationship between the amu mass and avogadro's number Avogadro's number is the amount of particles in 1 mole of a compound. Like it says in your book, 1 gram is avogadro's number in amu You could solve this with a crosstable 

OpenStudy (kaylala):

ok. i got it thanks @thomaster

thomaster (thomaster):

No problem :)

OpenStudy (kaylala):

@thomaster wait a minute. i'm mixed up how's you get 1.66x10^-24? when we're supposed to find how many grams is in 1 amu?

OpenStudy (kaylala):

@thomaster if the answer is 1.66x10^-24; then how did you get 1.66x10^-24 as your solution variable?

thomaster (thomaster):

1 divided by avogadro's number = 1.66*10^-24

thomaster (thomaster):

because 6.022*10^23 amu = 1 gram

OpenStudy (kaylala):

@thomaster where did 1 come from?

OpenStudy (kaylala):

what's the precise formula?

OpenStudy (kaylala):

what's the precise formula? and the summary of all? @thomaster

thomaster (thomaster):

It's explained here: 

thomaster (thomaster):

since 1*1 = 1 that's where the 1 comes from

OpenStudy (kaylala):

ok thanks again for clarifying :))))))